OpenStudy (shaleiah):

Find the inverse: f(x)=2x-1/3

satellite73 (satellite73):

assuming it is \[\frac{2x-1}{3}\], solve \[x=\frac{2y-1}{3}\]for \(y\)

satellite73 (satellite73):

any idea how to do that?

OpenStudy (shaleiah):

no :(

satellite73 (satellite73):

ok could you do it with numbers ? solve for \(y\) \[\frac{2y-1}{3}=7\] for example

OpenStudy (shaleiah):

y=11?

satellite73 (satellite73):

i don't know, i didn't do it but the first step is to multiply both sides by 3, to get \[2y-1=3\times 7\]

satellite73 (satellite73):

then add one, get \[2y=22\] then divide by 2, get \(y=11\) so you are right

satellite73 (satellite73):

now lets repeat with variables, the procedure is the same

OpenStudy (mathmale):

Would you please draw this equation, f(x)=2x-1/3, so that all involved here can be absolutely sure we're on the same wavelength?

satellite73 (satellite73):

\[\frac{2y-1}{3}=x\] first multiply both sides by \(3\)

OpenStudy (mathmale):

Note that satellite73 had to ask for clarification, and I too wondered what you meant.

OpenStudy (shaleiah):

this symbol "/" means division :) @mathmale

OpenStudy (mathmale):

As if I didn't know. But did you mean f(x)=2x-1/3 as in \[f(x)=2x-\frac{ 1 }{ 3 },\]

OpenStudy (mathmale):

or did you mean f(x)=2x-1/3 as in \[f(x)=(2x-1)/3, or f(x)=\frac{ 2x-1 }{ 3 }?\]

OpenStudy (shaleiah):

\[f(x)=\frac{ 2x-1 }{ 3}\]

OpenStudy (shaleiah):

2x-1=3x

OpenStudy (mathmale):

OK. Fine. In that case it is essential that y ou enclose "2x-1" inside parentheses, OR write your function as \[f(x)=\frac{ 2x-1 }{ 3 }\]

OpenStudy (mathmale):

2x-1=3x ? Where did that come from?

satellite73 (satellite73):

close \[2y-1=3x\] is step one

satellite73 (satellite73):

next add 1 to both sides

OpenStudy (shaleiah):

alright.

OpenStudy (mathmale):

Yes, but where did that come from? Please explain what you're doing.

OpenStudy (shaleiah):

2y-1+1=3x+1

OpenStudy (mathmale):

Find the inverse of f(x)=2x-1/3. Note that you MUST write this as f(x)=(2x-1)/3. Next, replace f(x) with y: y=(2x-1)/3. What is the next step, given that you are to find the inverse of f(x)?

satellite73 (satellite73):

did you get to \[2y=3x+1\] yet?

OpenStudy (shaleiah):

yes

satellite73 (satellite73):

last step, divide by 2 which really means just write it

OpenStudy (shaleiah):

\[\frac{ 3x+1 }{ 2 }\]

satellite73 (satellite73):

bingo \[f^{-1}(x)=\frac{3x+1}{2}\]

satellite73 (satellite73):

notice that \[f(11)=\frac{2\times 11-1}{3}=7\] and \[f^{-1}(7)=\frac{3\times 7+1}{2}=11\] just as it should

OpenStudy (shaleiah):

Got it :)